An automobiles velocity starting from rest is where v is measured in feet per second. Find
i) Its velocity after 5 seconds
ii) Its average velocity during the 10th second
iii) Its acceleration after 2 seconds
iv) Its average acceleration from t=2 to t=4
1
Expert's answer
2020-12-10T14:10:05-0500

i)Velocity after 5 sec is: V=V0+at,

V0=0, V=5a

ii) Average velocity Vavr= Stotal/ttotal

S=at2/2. In our case t=10 : Vavr=at2/(2×10), Vavr=5a

iii)acceleration after 2 sec, we can derive it from the formual:V=V0+at, from here we find a, a=V/t, a=V/2.

iv) average acceleration from t=2 to t=4 can be found by (a2+a4)/2.

a2=V2/2, a4=V4/4. So

aavr=(V2/2+V4/4)/2,

aavr=(2V2+V4)/8